OpenStudy (anonymous):

2. Suppose you drop a tennis ball from a height of 15 feet. After the ball hits the floor, it rebounds to 85% of its previous height. How high will the ball rebound after its third bounce? Round to the nearest tenth. a. Write an explicit formula for the sequence. Explain where you found the numbers you are putting in the formula. b. Identify the value of n and explain where you found it. Use the explicit formula to solve the problem. c. Write your final answer as a sentence.

OpenStudy (anonymous):

h(n+1) = (0.75) h(n) h(0) = 15 ft or h(n) = 15 (0.75)^n n=1 ftrst bounce, etc.